I am having trouble figuring out this problem.
The density of a penny is 8.93g/cm^3. the mass of a role of pennies is 177.6g. if the diameter of a penny is 1.85cm, how tall is the roll of pennies?
Please show me how to the solve the problem step by step.

Answer by mananth(16946)

Density = Mass / volume
.
D= 8.93g/cm^3
Mass = 177.6 g
Volume of the roll = m/d = 177.6 /9.83
Volume of roll = 8.06
= pi*0.925^2*H
.
H= 8.06*1/pi *1/(0.925)^2
= 2.998 cm
or 3cm is the height of the pile